OpenStudy (anonymous):

How to integrate 1/sin^2(x) Do I take u to be equal to sin^2(x)?

zepdrix (zepdrix):

\[\Large\rm \frac{1}{\sin x}=\csc x\]

zepdrix (zepdrix):

\[\Large\rm \int\limits \csc^2x ~dx\]So you can write it like this, yes?

zepdrix (zepdrix):

Do you remember your derivatives involving cosecant and cotangent and all those? :)

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

If I had to use integration by substitution what would u be? Will is be csc^2(x)

zepdrix (zepdrix):

I don't think you can do substitution. Hmm.

zepdrix (zepdrix):

You just need to remember that the derivative of \(\Large\rm -\cot x\) is \(\Large\rm \csc^2 x\).

OpenStudy (anonymous):

So the answer is -cot(x) +c?

zepdrix (zepdrix):

Mmmm yah that sounds right!